A supervisor records the repair cost for 14 randomly selected refrigerators. A sample mean of $79.20 and standa deviation of $10.41 are subsequently computed. Determine the % confidence interval for the mean repair cost for the refrigerators. Assume the population is approximately normal. Step 1 of 2: Find the critical value that should be used in constructing the confidence interval Round your answer to three decimal places Tables Keypad Answer How to enter your answer
Answer #1

Solution :

Given that,

t\alpha /2,df = 1.771

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 1.771 * (10.41 / \sqrt 14)

= 4.927

The 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

79.20 - 4.927 < \mu < 79.20 + 4.927

74.273 < \mu < 84.127

(74.273 , 84.127)

The 90% confidence interval for the mean : 74.273 , 84.127